# Learning to Instruct for Visual Instruction Tuning

### v1

@@ 3.1 Setup
Table 1: Performance comparison of VIT and LIT across 16 representative multimodal benchmarks. The benchmarks span four categories: General VQA, multi-discipline benchmarks, Chart/Document/OCR understanding, and image captioning. We report the average relative improvement for each category, along with the overall relative improvement across all benchmarks. *LLaVA-NeXT’s training script is not fully open-sourced; our implementation may differ slightly.
Benchmark Visual Instruction Tuning (VIT) Learning to Instruct (LIT)
@@
Metric Adv. Pop. Random $\text{CHAIR}_{s}^{\text{G}}$ ($\downarrow$) $\text{CHAIR}_{i}^{\text{G}}$ ($\downarrow$) $\text{CHAIR}_{s}^{\text{B}}$ ($\downarrow$) $\text{CHAIR}_{i}^{\text{B}}$ ($\downarrow$) $qAcc$ $fAcc$ $aAcc$
@@
LIT 2.36 0.53 3.08 1.67 2.92 1.92 2.25 3.33 1.92 1.83
3.2 Main Results
In this section, we present a comprehensive evaluation of LIT across multiple benchmarks, as shown in Table 1.
General Visual Question Answering. We first evaluate several widely used VQA benchmarks, including the most prominent VQAv2 benchmark and the visual reasoning-oriented GQA. We also assess VizWiz, which requires predicting unanswerable questions, and ScienceQA, which covers a diverse range of scientific topics. Our results demonstrate that our proposed LIT achieves competitive VQA performance when compared to VIT.
Comprehensive Multimodal Benchmarks. We evaluate the multi-discipline multimodal understanding and reasoning capabilities of our LIT through MME, MMMU and MMStar benchmarks. These benchmarks span a wide range of disciplines, including Art & Design, Business, Science, Health & Medicine, Humanities & Social Sciences, Technology & Engineering, and encompass diverse tasks such as object presence, counting, spatial reasoning, commonsense inference, numerical computation, and code reasoning. Our results demonstrate that LIT consistently outperforms VIT, achieving an average relative improvement of up to 3.7%. This underscores the effectiveness of LIT in enhancing performance across diverse multimodal applications.
Chart, Document, and OCR Understanding. We include a diverse set of OCR-related datasets to evaluate the performance of models in locating, extracting and interpreting text from various data visualizations, including charts, diagrams, documents, and natural images, in order to accurately address the associated questions. The results indicate that our LIT significantly outperforms VIT across all OCR-related benchmarks, yielding an average relative improvement of 6.3%. This highlights the effectiveness of the proposed LIT in extracting fine-grained visual information. This can be attributed to the fact that our LIT encourages the proactive generation of instructions. In OCR-related datasets like TextVQA, the instructions require predicting “Reference OCR token” that captures the informative textual content within the image. This necessitates a more comprehensive understanding of the low-level information embedded in the images, thus enhancing OCR capabilities.
Image Captioning. We assess image-captioning, a representative image-text task widely used as a pretraining task for MLLMs. Our analysis includes classical captioning datasets such as COCO2017 and Flickr30k, along with more specialized datasets such as TextCaps, which involves fine-grained textual content description, and NoCaps, which represents out-of-domain scenarios. The results demonstrate that LIT achieves the most substantial performance improvement over VIT on captioning datasets, among the various capabilities evaluated, with an average relative improvement of up to 17.6%. The reasons are two-fold: (1) LIT motivates a more comprehensive understanding of the images, which guarantees the generation of accurate descriptions, and (2) LIT alleviates the negative effects of overfitting on instruction-following abilities, preserving more of the captioning skills acquired during the pretraining stage.
Finding. LIT consistently outperforms VIT across all benchmarks, achieving an average relative improvement of up to 8.5%. The most significant gains are observed in OCR-related and captioning datasets, with improvements of up to 8.8% and 17.6%, respectively. This suggests that LIT excels at regularizing MLLMs to pay more attention to visual input and strengthening their fundamental visual capabilities.
3.3 Hallucination Evaluation
In this section, we evaluate the effectiveness of our method in mitigating hallucinations, as shown in Table 2.
@@ 3.3 Hallucination Evaluation
Results on POPE. POPE (Li et al., 2023b) is the most commonly used benchmark for object hallucination. It consists of “Yes or No” questions to evaluate the MLLMs’ capability to determine whether the given object is in the image. We compute the accuracy based on questions derived from adeversarial, popular and random sampling. Our results demonstrate that LIT outperforms VIT in both adversarial and popular sampling settings, showcasing the effectiveness of LIT in reducing object hallucinations.
@@
Results on CHAIR Evaluation. In addition to the discriminative evaluations on POPE, we utilize the CHAIR metric (Rohrbach et al., 2018) as a complementary approach to evaluate object hallucination in image captioning tasks. Specifically, CHAIR quantifies the proportion of objects referenced in an image caption that are absent from the corresponding ground-truth label set. As shown in Table 2, our LIT consistently outperforms VIT, achieving improvements of 2.2% and 1.7% in terms of the $\text{CHAIR}_{s}$ and $\text{CHAIR}_{i}$ metric, respectively. This further demonstrates that our LIT effectively regularize MLLMs, reducing the generation of plausible but visually irrelevant contents.
(a) Training loss on VQAv2.
@@
Figure 7: (a) Training loss distribution of LIT and VIT on the VQAv2 dataset. (b) Testing loss distribution of LIT and VIT on the DocVQA dataset. The standard cross-entropy loss is calculated for the response part of each data sample. Experiments are based on TinyLLaVA Qwen2-0.5B.
@@
Figure 8: Ablation on (a) instruction data scale and (b) pretraining data scale. The figure shows the performance of LIT and VIT using 40%, 60%, 80%, and 100% of the data, with another stage using the full dataset. Experiments are based on TinyLLaVA Qwen2-0.5B.
Results on MMHAL-Bench. CHAIR mainly focuses on object hallucination by evaluating the presence of objects. To provide a more fine-grained assessment, we further adopt MMHAL-Bench (Sun et al., 2024a) to evaluate a broader range of hallucination types. For evaluating the quality of the generated content, we leverage GPT-4 to compute the GPT score and hallucination rate for comparison. The results demonstrate that LIT significantly outperforms VIT by 36% in average score and 22% in hallucination rate. Moreover, LIT consistently outperforms VIT across various hallucination types, with notable improvements in adversarial, comparison, relation, environment, and holistic settings.
Results on HallusionBench. To further improve the comprehensiveness of hallucination evaluation, we use HallusionBench (Guan et al., 2024) to include more scenarios with various disciplines, image types and visual input modalities. We report Question Pair Accuracy ($qAcc$), Figure Accuracy ($fAcc$) and All Accuracy ($aAcc$) for hallucination assessment based on GPT-4 for comparison. The results demonstrate that LIT outperforms VIT, with improvements of 0.44% and 1.05% in terms of $qAcc$ and $aAcc$, respectively, highlighting its effectiveness in mitigating a wider range of failure modes in multimodal hallucinations.
Table 3: Ablations on template removal. The table shows the performance of VIT, which learns to respond on $X_{A}$, and LIT, which learns to instruct using variants of progressively removing system template $\mathbf{X}_{I}^{S}$ and task template $\mathbf{X}_{I}^{T}$. Here, $X_{I}^{\backslash S,T}$ refers to the instructions excluding both the system and task templates. Experiments are based on TinyLLaVA Qwen2-0.5B.
@@
$X_{I}^{\backslash S,T}$ $\mathbf{X}_{I}^{T}$ $\mathbf{X}_{I}^{S}$ $X_{A}$ VQAv2-L VizWiz-L MMMU MMStar ChartQA OCR Bench Flickr30k-L RefCOCO-L
@@
Table 4: Performance comparison of VIT and LIT when using a single data type as the primary training data, including Choice, Grounding, GPT-generated, Captioning, and QA data. Experiments are based on TinyLLaVA Qwen2-0.5B.
Data Method VQAv2-L MMMU ChartQA RefCOCO-L $\Delta$
Choice VIT 52.76 31.56 12.56 20.45 -
LIT 50.90 31.22 11.88 25.33 +3%
Grounding VIT 53.10 33.76 11.32 25.01 -
LIT 52.62 32.11 11.32 26.51 +0%
GPT-gen VIT 53.90 31.11 11.60 8.42 -
LIT 55.16 31.78 11.68 11.34 +10%
Captioning VIT 45.12 29.44 11.52 16.02 -
LIT 47.46 31.11 12.00 16.85 +5%
QA VIT 63.26 31.00 11.84 16.94 -
LIT 62.28 33.00 12.36 29.69 +21%
3.4 Further Analysis
Mitigate Overfitting from the Loss Perspective. In Figure 7, we present the loss distributions of our LIT and VIT on the training dataset (VQAv2) and test datasets (DocVQA and NoCaps). The cross-entropy loss is calculated specifically for the response component, excluding the instruction part, even for LIT. The results reveal that LIT exhibits slightly higher loss than VIT on the training data but achieves lower loss on unseen test data. This clearly demonstrates that our method, by learning to instruct images as a regularizer, effectively avoids overfitting on the training data and exhibits superior generalization performance.
Ablation on the Scale of Instruction Data. To investigate the impact of different data volumes during the instruction tuning phase on our method, we present the performance of both the baseline VIT and our LIT at 40%, 60%, 80% data levels in Figure 8(a). Note that we use the same scale for the axes of both radar charts. It can be observed that, at the same data scale, our method significantly outperforms VIT. Specifically, from 40% to 80% instruction data, our method shows an average performance improvement over VIT across the eight metrics in Figure 8(a) by 5.9%, 13.0%, and 9.3%, respectively. The performance improvements suggest that our method benefits from increasing data volume, highlighting its potential for continuous improvement within the research trend of expanding data scaling laws.
Ablation on the Scale of Pretraining Data. We also investigate the impact of different data volumes during the pretraining phase on our method, with the results shown in Figure 8(b). Our LIT demonstrates a clear advantage over the baseline at the same data scale. From 40% to 80% pretraining data, the average improvements are 10.9%, 8.7%, and 8.5%, respectively. Interestingly, the performance with 80% pretraining data even surpasses that with 100% pretraining data. This “less-is-more” (Zhou et al., 2023) phenomenon highlights the complexity of how pretraining data impacts final performance.
Ablation on Types of instruction data. We selected different types of data from the full dataset (mixed with 10% of other types of data), including QA data, GPT-generated data, selection data, grounding data, and captioning data.We conduct instruction tuning using each of these data types, and the results are presented in Table 4. It can be observed that the improvements of our method vary considerably across different data types, which we attribute primarily to the differences in instruction quality. The instructions for grounding data consist only of fixed templates and almost random bounding box coordinates, making it difficult for the model to gain improvements from learning such instructions. In contrast, the instructions for QA data contain a wealth of information related to the image content, which enables our method to provide a significant improvement.
Ablation on Template Removal. To gain more insights, we conduct experiments to assess the impact of removing each prompt template from the instructions, as shown in Table 3. Four experimental setups are considered: (1) the baseline method, which learns only the answer component; (2) the straightforward method, which learns both the full set of instructions and answers; (3) the method that learns the instructions excluding system and format messages (USER/ASSISTANT tokens), while still learning the answers; and (4) the method that learns the instructions without system/format messages and task-specific prompts, while learning the answers as well. The results indicate that progressively removing task-irrelevant template tokens enhances the overall performance of LIT, preventing MLLMs from overfitting to redundant template tokens.
Computational Analysis. To validate the computational efficiency of LIT, we conduct experiments on finetuning LLaVA-1.5 Vicuna-7B using instructional data with varying instruction-to-response length ratios ($L_{I}/L_{A}$), ranging from 0.05 to 20. The data are sampled from the training set. We report the average number of samples per second and the average number of steps per second (averaged over 100 training steps). Figure 9 demonstrates that LIT achieves an average of $0.331\pm 0.005$ steps per second, while VIT achieves $0.334\pm 0.005$. Our LIT incur only a negligible computational overhead of less than 1%. Detailed results can be referred to Table 8 in the Appendix E.

### v2

@@ 3.2 Main Results
Table 1: Performance comparison of VIT and L2T across 16 representative multimodal benchmarks. The benchmarks span four categories: General VQA, multi-discipline benchmarks, Chart/Document/OCR understanding, and image captioning. We report the average relative improvement for each category, along with the overall relative improvement across all benchmarks. *LLaVA-NeXT’s training script is not fully open-sourced; our implementation may differ slightly.
Benchmark Visual Instruction Tuning (VIT) Learning to Instruct (L2T)
@@
General Visual Question Answering. We first evaluate several widely used VQA benchmarks, including the most prominent VQAv2 benchmark and the visual reasoning-oriented GQA. We also assess VizWiz, which requires predicting unanswerable questions, and ScienceQA, which covers a diverse range of scientific topics. Our results demonstrate that our proposed L2T achieves competitive VQA performance when compared to VIT.
Comprehensive Multimodal Benchmarks. We evaluate the multi-discipline multimodal understanding and reasoning capabilities of our L2T through MME, MMMU and MMStar benchmarks. These benchmarks span a wide range of disciplines, including Art & Design, Business, Science, Health & Medicine, Humanities & Social Sciences, Technology & Engineering, and encompass diverse tasks such as object presence, counting, spatial reasoning, commonsense inference, numerical computation, and code reasoning. Our results demonstrate that L2T consistently outperforms VIT, achieving an average relative improvement of up to 3.7%. This underscores the effectiveness of L2T in enhancing performance across diverse multimodal applications.
Chart, Document, and OCR Understanding. We include a diverse set of OCR-related datasets to evaluate the performance of models in locating, extracting and interpreting text from various data visualizations, including charts, diagrams, documents, and natural images, in order to accurately address the associated questions. The results indicate that our L2T significantly outperforms VIT across all OCR-related benchmarks, yielding an average relative improvement of 6.3%. This highlights the effectiveness of the proposed L2T in extracting fine-grained visual information. This can be attributed to the fact that our L2T encourages the proactive generation of instructions. In OCR-related datasets like TextVQA, the instructions require predicting “Reference OCR token" that captures the informative textual content within the image. This necessitates a more comprehensive understanding of the low-level information embedded in the images, thus enhancing OCR capabilities.
Image Captioning. We assess image-captioning, a representative image-text task widely used as a pretraining task for MLLMs. Our analysis includes classical captioning datasets such as COCO2017 and Flickr30k, along with more specialized datasets such as TextCaps, which involves fine-grained textual content description, and NoCaps, which represents out-of-domain scenarios. The results demonstrate that L2T achieves the most substantial performance improvement over VIT on captioning datasets, among the various capabilities evaluated, with an average relative improvement of up to 17.6%. The reasons are two-fold: (1) L2T motivates a more comprehensive understanding of the images, which guarantees the generation of accurate descriptions, and (2) L2T alleviates the negative effects of overfitting on instruction-following abilities, preserving more of the captioning skills acquired during the pretraining stage.
Finding. L2T consistently outperforms VIT across all benchmarks, achieving an average relative improvement of up to 8.5%. The most significant gains are observed in OCR-related and captioning datasets, with improvements of up to 8.8% and 17.6%. This suggests that L2T excels at regularizing MLLMs to pay more attention to visual input and strengthening their fundamental visual capabilities.
3.3 Hallucination Evaluation
In this section, we evaluate the effectiveness of L2T in mitigating hallucinations, as shown in Table˜2.
@@ 3.3 Hallucination Evaluation
Metric Adv. Pop. Random $\text{CHAIR}_{s}^{\text{G}}$ ($\downarrow$) $\text{CHAIR}_{i}^{\text{G}}$ ($\downarrow$) $\text{CHAIR}_{s}^{\text{B}}$ ($\downarrow$) $\text{CHAIR}_{i}^{\text{B}}$ ($\downarrow$) qAccqAcc fAccfAcc aAccaAcc
@@
L2T 2.36 0.53 3.08 1.67 2.92 1.92 2.25 3.33 1.92 1.83
@@
Results on POPE. POPE (Li et al., 2023b) is the most commonly used benchmark for object hallucination. It consists of “Yes or No" questions to evaluate the MLLMs’ capability to determine whether the given object is in the image. We report accuracies based on questions derived from adeversarial, popular and random sampling. L2T outperforms VIT in both adversarial and popular sampling settings, showcasing the effectiveness of L2T in reducing object hallucinations.
@@
Results on CHAIR Evaluation. In addition to the discriminative evaluations on POPE, we utilize the CHAIR metric (Rohrbach et al., 2018) as a complementary approach to evaluate object hallucination in image captioning tasks. Specifically, CHAIR quantifies the proportion of objects referenced in an image caption that are absent from the corresponding ground-truth label set. As shown in Table˜2, our L2T consistently outperforms VIT, achieving improvements of 2.2% and 1.7% in terms of the $\text{CHAIR}_{s}$ and $\text{CHAIR}_{i}$ metric, respectively. This further demonstrates that our L2T effectively regularize MLLMs, reducing the generation of plausible but visually irrelevant contents.
Results on MMHAL-Bench. CHAIR mainly focuses on object hallucination by evaluating the presence of objects. To provide a more fine-grained assessment, we further adopt MMHAL-Bench (Sun et al., 2024b) to evaluate a broader range of hallucination types. For evaluating the quality of the generated content, we leverage GPT-4 to compute the GPT score and hallucination rate for comparison. The results demonstrate that L2T significantly outperforms VIT by 36% in average score and 22% in hallucination rate. Moreover, L2T consistently outperforms VIT across various hallucination types, with notable improvements in adversarial, comparison, relation, environment, and holistic settings.
Results on HallusionBench. To further improve the comprehensiveness of hallucination evaluation, we use HallusionBench (Guan et al., 2024) to include more scenarios with various disciplines, image types and input modalities. We report Question Pair Accuracy ($qAcc$), Figure Accuracy ($fAcc$) and All Accuracy ($aAcc$) based on GPT-4 for comparison. The results demonstrate that L2T outperforms VIT, with improvements of 0.44% and 1.05% in terms of $qAcc$ and $aAcc$, respectively, highlighting its effectiveness in mitigating a wider range of failure modes in multimodal hallucinations.
3.4 Further Analysis
(a) Train loss on VQAv2.
@@ 3.4 Further Analysis
Figure 8: (a) Training loss distribution of L2T and VIT on VQAv2. (b) Testing loss distribution of L2T and VIT on DocVQA. The standard cross-entropy loss is calculated for the response part. Experiments are based on TinyLLaVA Qwen2-0.5B.
@@
Figure 9: Ablation on (a) instruction data scale and (b) pretraining data scale. The figure shows the performance of L2T and VIT using 40%, 60%, 80%, and 100% of the data, with another stage using the full dataset. Experiments are based on TinyLLaVA Qwen2-0.5B.
Mitigate Overfitting from the Loss Perspective. In Figure˜9, we present the loss distributions of our L2T and VIT on the training dataset (VQAv2) and test datasets (DocVQA and NoCaps). The cross-entropy loss is calculated specifically for the response component, excluding the instruction part, even for L2T. The results reveal that L2T exhibits slightly higher loss than VIT on the training data but achieves lower loss on unseen test data. This clearly demonstrates that our method, by learning to instruct images as a regularizer, effectively avoids overfitting on the training data and exhibits superior generalization performance.
Ablation on the Scale of Instruction Data. To investigate the impact of different data volumes during the instruction tuning phase on L2T, we present the performance of both the baseline VIT and L2T at 40%, 60%, 80% data levels in Figure˜9(a). Note that we use the same scale for the axes of both radar charts. It can be observed that, at the same data scale, L2T significantly outperforms VIT. Specifically, from 40% to 80% instruction data, our method shows an average performance improvement over VIT across the eight metrics in Figure˜9(a) by 5.9%, 13.0%, and 9.3%, respectively. The performance improvements suggest that our method benefits from increasing data volume, highlighting its potential for continuous improvement within the research trend of expanding data scaling laws.
Ablation on the Scale of Pretraining Data. We also investigate the impact of data volumes during the pretraining phase on L2T, with the results shown in Figure˜9(b). L2T demonstrates a clear advantage over the baseline at the same data scale. From 40% to 80% pretraining data, the average improvements are 10.9%, 8.7%, and 8.5%, respectively. Interestingly, the performance with 80% pretraining data even surpasses that with 100% pretraining data. This “less-is-more" (Zhou et al., 2023) phenomenon highlights the complexity of how pretraining data impacts final performance.
Table 3: Performance comparison of VIT and L2T when using a single data type as the primary training data, including Choice, Grounding, GPT-generated, Captioning, and QA data. Experiments are based on TinyLLaVA Qwen2-0.5B.
Data Method VQAv2-L MMMU ChartQA RefCOCO-L $\Delta$
Choice VIT 52.76 31.56 12.56 20.45 -
L2T 50.90 31.22 11.88 25.33 +3%
Grounding VIT 53.10 33.76 11.32 25.01 -
L2T 52.62 32.11 11.32 26.51 +0%
GPT-gen VIT 53.90 31.11 11.60 8.42 -
L2T 55.16 31.78 11.68 11.34 +10%
Captioning VIT 45.12 29.44 11.52 16.02 -
L2T 47.46 31.11 12.00 16.85 +5%
QA VIT 63.26 31.00 11.84 16.94 -
L2T 62.28 33.00 12.36 29.69 +21%
Ablation on Types of instruction data. We selected different types of data from the full dataset (mixed with 10% of other types of data), including QA data, GPT-generated data, selection data, grounding data, and captioning data.We conduct instruction tuning using each of these data types, and the results are presented in Table˜3. It can be observed that the improvements of our method vary considerably across different data types, which we attribute primarily to the differences in instruction quality. The instructions for grounding data consist only of fixed templates and almost random bounding box coordinates, making it difficult for the model to gain improvements from learning such instructions. In contrast, the instructions for QA data contain a wealth of information related to the image content, which enables our method to provide a significant improvement.
[…]
@@
[…]
@@
Ablation on Template Removal. To gain more insights, we conduct experiments to assess the impact of removing each prompt template from the instructions, as shown in Table˜4. Four experimental setups are considered: (1) the baseline method, which learns only the answer component; (2) the straightforward method, which learns both the full set of instructions and answers; (3) the method that learns the instructions excluding system and format messages (USER/ASSISTANT tokens), while still learning the answers; and (4) the method that learns the instructions without system/format messages and task-specific prompts, while learning the answers as well. The results indicate that progressively removing task-irrelevant template tokens enhances the overall performance of L2T, preventing MLLMs from overfitting to redundant template tokens.
@@
Computational Analysis. To validate the computational efficiency of L2T, we conduct experiments on finetuning LLaVA-1.5 Vicuna-7B using instructional data with varying instruction-to-response length ratios ($L_{I}/L_{A}$), ranging from 0.05 to 20. The data are sampled from the training set. We report the average number of samples per second and the average number of steps per second (averaged over 100 training steps). Figure˜10 demonstrates that L2T achieves an average of $0.331\pm 0.005$ steps per second, while VIT achieves $0.334\pm 0.005$. Our L2T incur only a negligible computational overhead of less than 1%. Detailed results can be referred to Table˜11 in the Appendix˜E.
Evaluation on another VLM baseline Prism-7B. To further verify the generalizability of L2T, we conduct additional experiments on another robust vision-language model (VLM) baseline, Prism (Karamcheti et al., 2024), which surpasses LLaVA-1.5 by incorporating optimized training strategies, advanced image preprocessing, and fused visual backbones such as SigLIP and DiNOv2. We apply L2T to Prism-DINOSigLIP-Controlled-7B and compare it with the standard VIT-style Prism baseline across diverse benchmarks, including visual question answering (GQA, VizWiz, TextVQA), localization (RefCOCO, RefCOCO+, RefCOCOg), and challenging reasoning tasks (POPE, VSR, AI2D). As shown in Table˜6, L2T consistently improves performance across all benchmarks, demonstrating its effectiveness and broad applicability.
Table 5: Performance comparison on text-only benchmarks. L2T maintains comparable or improved performance, indicating no degradation in core language ability.
Benchmark MTBench WildBench MMLU AGIEval
ViT LLaVA1.5-7B 5.35 -9.40 49.57 32.62
LiT LLaVA1.5-7B 5.49 -6.27 49.18 32.49
Effect on Textual Understanding. To examine whether L2T compromises language proficiency by altering the SFT data composition, we evaluate the model on several text-only benchmarks, including MT-Bench, WildBench, MMLU, and AGIEval. We use the LLaVA-1.5 Vicuna-7B model as the baseline and apply L2T under identical training configurations. As shown in Table˜5, L2T achieves comparable results to the baseline on MT-Bench, MMLU, and AGIEval, while notably improving WildBench performance. These results indicate that learning to instruct functions as a synergistic regularization rather than a disruptive modification, enhancing visual grounding without compromising the model’s core text comprehension capabilities.
Table 6: Performance comparison on VLM baseline Prism-7B. L2T yields consistent improvements across diverse benchmarks, highlighting its general effectiveness.
Task GQA VizWiz TextVQA RefCOCO RefCOCO+ RefCOCOg POPE VSR AI2D
VIT Prism-7B 61.92 55.36 52.80 56.70 50.70 52.70 88.00 53.20 55.50
L2T Prism-7B 62.62 57.75 55.60 66.00 58.90 62.00 88.50 61.70 57.10
Exploring the potential of L2T for self-improving instruction tuning. To illustrate the broader potential of L2T, we conduct a pilot exploration where the model leverages its own generation capability to enhance itself. Starting from a L2T model pretrained on a 100k subset of the LLaVA-mix-665k dataset, we prompt it with image-only inputs to automatically produce 100k instruction–response pairs. These self-generated samples are then merged with the original training data for continued fine-tuning. As shown in Table˜7, this simple self-bootstrapping process leads to consistent gains across multiple benchmarks. The results highlight L2T’s potential to evolve through its own generated supervision, paving the way toward self-improving vision–language models.
Table 7: Exploring the potential of L2T for self-improving instruction tuning. Incorporating 100k model-generated instruction–response pairs improves performance across diverse benchmarks.
Benchmark VQAv2-L VizWiz-L MMMU MMSTAR ChartQA OCR Bench Flickr30k-L RefCOCO-L
L2T TinyLLaVA-0.5B 56.68 23.20 31.78 33.46 11.96 23.40 68.55 17.42
L2T TinyLLaVA-0.5B w/ generated data 60.80 24.22 32.33 34.75 12.44 24.50 68.32 24.53